OpenStudy (anonymous):

What is the slope intercept form of the equation 3x-y= - 6? What is the slope intercept form of the equation x-3y=3? What is the slope intercept form of the equation 9x+3y=6? What is the slope intercept form of the equation x+3y=-6

OpenStudy (anonymous):

a = y = -1 ----x - 2 3 b= y = 1 --- x - 1 3 c = y - 3x+2 d = y -3x + 6

OpenStudy (anonymous):

Subtract 3x from both the sides and tell me what you got..

OpenStudy (anonymous):

x = y/3 - 2

OpenStudy (anonymous):

I told you to subtract and not to divide etc: \[3x - y - 3x = -6 - 3x\] Tell me after this what you will get..

OpenStudy (anonymous):

3x